Dwight Donovan, the president of Benson Enterprises, is considering two investment opportunities. Because of limited resources, he will be able to invest in only one of them. Project A is to purchase a machine that will enable factory automation; the machine is expected to have a useful life of five years and no salvage value. Project B supports a training program that will improve the skills of employees operating the current equipment. Initial cash expenditures for Project A are $116,000 and for Project B are $36,000. The annual expected cash inflows are $29,823 for Project A and $10,486 for Project B. Both investments are expected to provide cash flow benefits for the next five years. Benson Enterprises’ cost of capital is 8 percent. (PV of $1 and PVA of $1) (Use appropriate factor(s) from the tables provided.)

Required

  1. Compute the net present value of each project. Which project should be adopted based on the net present value approach?

  2. Compute the approximate internal rate of return of each project. Which one should be adopted based on the internal rate of return approach?

net present value of project A and project B?

internal rate of return of project a and project B?
Answer #1
a] NPV of Project A = 29823*3.9927-116000 = $          3,074.29
NPV of Project B = 10486*3.9927-36000 = $          5,867.45
Project B should be adopted as it has higher NPV.
b] IRR is that discount rate for which NPV = 0.
So, for Project A
0 = 29823*PVIFA(irr,5)-116000
PVIFA(irr,5) = 116000/29823 = 3.8896
From the PVIFA tables, the factor for 9% for
5 years is 3.8897. Hence, IRR estimate for
Project A = 9.00%
Similarly, for Project B
PVIFA(irr,5) = 36000/10486 = 3.4331
From the PVIFA tables, the factor for 14% for
5 years is 3.4331. Hence, IRR for Project B = 14.00%
Project B should be adopted as it has higher IRR.